If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are ________?
Answer by checkley75(3666)

72/3=24 MPH DOWNSTREAM
60/6=10 MPH UP STREAM
(24-10)/2=14/2=7 MPH FOR THE CURRENT &
24-7=17 MPH FOR THE BOAT OR 10+7=17 MPH FOR THE BOAT